In the game of blackjack, the player is initially dealt two cards from a deck of ordinary playing cards. Without going into all the details of the game, it will suffice to say here that the best possible hand one could receive on the initial dear is a combination of an ace of any suit and any face card or 10. What is the probability that the player will be dealt this combination?
Answer #1

First we need to find out the no of combinations of one ace and any one card out of 10, J, Q and K out of the 52 cards.

In total there are 4 aces and 4 x 4 = 16 total cards among 10, J, Q and K.

So total no of such combinations = 4 x 16 = 64

Now any 2 cards out of 52 cards can be selected in 52C2 = 1326

So the required probability = 64/1326 = 32/663 = 0.0483